OpenStudy (anonymous):

Find the indicated sum. 6+9+12+15+⋯+183

OpenStudy (anonymous):

this is an arithmetic progression with the 1st term A = 6, common difference D = 3, and the last term L = 183. We need to find the no. of terms N now L = A + (N-1)D 183 = 6 + (N-1)3 177 = (N-1)3 59 = N-1 OR N = 60 Now sum of an A.P. = N/2 ( A+L) A = FIRST TERM OF SERIES; L = LAST TERM OF SERIES => 60/2(6+183) => 30 X 189 = 5670 is the answer. :)

OpenStudy (anonymous):

What is first term and common difference here??

OpenStudy (anonymous):

6 and 3

OpenStudy (anonymous):

Yep.. And you have last term an given, so : \[a_n = a + (n -1)d \implies 183 = 6 + (n - 1) \cdot 3\] Can you find n here??

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Yep.. So you are one step away now from answer: Use: \[\large \color{green}{Sum = \frac{n}{2}(a + a_n)}\]

OpenStudy (anonymous):

i have the answer already

OpenStudy (anonymous):

Okay, then best of luck...